OpenStudy (anonymous):

Finding the ARea of the Region f(x)= 2sinx + sin2x y=0

OpenStudy (anonymous):

\[0 \le x \le \pi\] interval

OpenStudy (anonymous):

I have this so far \[\int\limits_{0}^{\pi} (2\sin x + \sin2x) dx\]

OpenStudy (anonymous):

I have to integrate...so I got this... \[-2\cos x + ??? \]

OpenStudy (xapproachesinfinity):

seems to me you did a good start

OpenStudy (anonymous):

I don't get how to integrate sin2x ? i Know sinx = -cosx but what if it has 2 in there?

OpenStudy (xapproachesinfinity):

\[\int_{0}^{\pi}2\sin x dx+\int_{0}^{\pi}\sin 2x dx=-2\cos|_{0}^{\pi}+\int2\sin x\cos xdx\]

OpenStudy (xapproachesinfinity):

can you intgerate that sin 2x now

OpenStudy (xapproachesinfinity):

the first one is easy we are done now focus on int (sin 2x ) which is the same as 2int(cosx sinx )dx

OpenStudy (anonymous):

oh okay...

OpenStudy (anonymous):

so does the 2 come to the front of sin ? will it be -2cosx ?

jimthompson5910 (jim_thompson5910):

xapproachesinfinity used the identity sin(2x) = 2*sin(x)*cos(x)

OpenStudy (xapproachesinfinity):

yes! exactly like jim said

jimthompson5910 (jim_thompson5910):

although now that I look at it, it's not a good move to use it

OpenStudy (anonymous):

whoops :/

jimthompson5910 (jim_thompson5910):

use u-substitution let u = 2x

OpenStudy (xapproachesinfinity):

you can also use substitution either way work

jimthompson5910 (jim_thompson5910):

oh true, you can

OpenStudy (anonymous):

i like the u substitution

jimthompson5910 (jim_thompson5910):

if you let u = 2x, then du = 2*dx ---> dx = du/2

OpenStudy (anonymous):

i have 1/2 = du

OpenStudy (anonymous):

i am not too great at trig sorry @xapproachesinfinity

OpenStudy (xapproachesinfinity):

replace dx be du/2 so to get int(sinudu/2)

OpenStudy (anonymous):

yep!! but then the 1/2 goes infront of the integral right ?

OpenStudy (xapproachesinfinity):

\[\int_{0}^{\pi}\sin u\frac{du}{2}\] 1/2 is constant so you can factor it out \[\frac{1}{2}\int_{0}^{\pi}\sin u du\]

OpenStudy (xapproachesinfinity):

yes correct 1/2 goes in front :)

OpenStudy (xapproachesinfinity):

oh the limits change though

jimthompson5910 (jim_thompson5910):

xapproachesinfinity, the limits aren't going to be 0 and pi if u = 2x, then when x = 0 ---> u = 2x = 2(0) = 0 when x = pi ---> u = 2(pi) = 2pi so the limits are now from 0 to 2pi

OpenStudy (xapproachesinfinity):

not 0 to pi

jimthompson5910 (jim_thompson5910):

yeah you caught it as I was writing

OpenStudy (xapproachesinfinity):

yea i just caught that :) thanks

jimthompson5910 (jim_thompson5910):

or you can just revert back to x (by plugging u = 2x back in) and not worry about changing the limits

jimthompson5910 (jim_thompson5910):

after you integrate

OpenStudy (xapproachesinfinity):

yea that's another way to do it well as long as the asker is with us lol

OpenStudy (xapproachesinfinity):

did you understand the u sub we just did?

OpenStudy (anonymous):

yes but i'm stuck on putting the u and du back in now before I integrate

OpenStudy (xapproachesinfinity):

all you need to do now is integrate \[1/2\int\sin udu\]

OpenStudy (xapproachesinfinity):

forgot the limits for now we gonna go back to it

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\int\limits_{0}^{\pi} du * u ??\]

OpenStudy (xapproachesinfinity):

no where did you get u here is what we did let u=2x du=2dx so dx=1/2 du \[\int\sin2xdx=\int\sin( u)\frac{du}{2}=1/2\int\sin( u)du\]

OpenStudy (xapproachesinfinity):

just repeated the us sub so to make sure you got it

OpenStudy (xapproachesinfinity):

now it is an easy integral just think what function gives sin x when differentiate it?

OpenStudy (anonymous):

ok here is what i have \[\int\limits_{0}^{\pi} \sin(u) du\]

OpenStudy (anonymous):

o h crap i forgot the 1/2

OpenStudy (xapproachesinfinity):

no not 0 to pi forgot those limits for now just integrate

OpenStudy (xapproachesinfinity):

just find the anti derivative first

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } \int\limits_{0}^{\pi} \sin(u) du\]

OpenStudy (xapproachesinfinity):

but not 0 to pi okay that's why i said forgot the limits for now

OpenStudy (xapproachesinfinity):

that's an error in my side as jim mentioned

OpenStudy (anonymous):

okay so are we now just finding the antiderv of sin2x right ? or the whole thing ?

OpenStudy (xapproachesinfinity):

we find the antiderivative of sin u which will give us the antiderivative of sin 2x

OpenStudy (xapproachesinfinity):

not the whole think of course one part is done already

OpenStudy (xapproachesinfinity):

thing*

OpenStudy (anonymous):

oh okay!! i got \[\frac{ 1 }{ 2 } -\cos2x\]

OpenStudy (anonymous):

can it be -1/2cos2x ?

OpenStudy (xapproachesinfinity):

okay but you forgot 1/2 that was attached to it

OpenStudy (xapproachesinfinity):

that is good but you need to multiply by 1/2

OpenStudy (anonymous):

-1/2cosx is that right ?

OpenStudy (xapproachesinfinity):

\[1/2\int \sin udu=-1/2\cos u=-1/2\cos (2x)\] sorry you got the correct one

OpenStudy (xapproachesinfinity):

no the whole thing \[=-2\cos x|_{0}^{\pi}+(-1/2 \cos 2x|_{0}^{\pi}\]

OpenStudy (xapproachesinfinity):

just evaluate now

OpenStudy (xapproachesinfinity):

sorry my attention is little spread heheh

OpenStudy (anonymous):

okay im going to evaluate

OpenStudy (xapproachesinfinity):

ok

OpenStudy (anonymous):

-2cosx - 1/2cosx

OpenStudy (anonymous):

2-1/2 = 3/2

OpenStudy (anonymous):

i plugged in pi

OpenStudy (xapproachesinfinity):

hmm you didn't do it right \[(-2\cos\pi+2\cos 0)+(-1/2 \cos 2\pi+1/2\cos0)\] is this what you did

OpenStudy (xapproachesinfinity):

you should have 2+2-1/2 +1/2=4

OpenStudy (anonymous):

oh no i didn't do that...but now i will !

OpenStudy (xapproachesinfinity):

okay

OpenStudy (xapproachesinfinity):

well how do you evaluate the integral \[\int _{a}^{b}f(x)dx=(F(x)|_{a}^{b}=F(b)-F(a)\]

OpenStudy (anonymous):

i got zero ?? what's going on :(

OpenStudy (xapproachesinfinity):

well check again

OpenStudy (anonymous):

okay i see, I subtracted instead of adding

OpenStudy (xapproachesinfinity):

what did you get

OpenStudy (anonymous):

still 0

OpenStudy (xapproachesinfinity):

hahaha show me your work

OpenStudy (anonymous):

this is what i have -2cos(0) + 1/2cos(0) + -2cos(pi) +1/2cos(pi)

OpenStudy (xapproachesinfinity):

hmm are you sure check what i wrote?

OpenStudy (xapproachesinfinity):

where is 2pi in there

OpenStudy (anonymous):

now i see the error!! okay....here goes

OpenStudy (anonymous):

why did we change the interval? i know it happened earlier, but why ?

OpenStudy (xapproachesinfinity):

http://www.wolframalpha.com/input/?i=int_%7B0%7D%5E%7Bpi%7D%282sinx%2Bsin2x%29dx check here the same result i got

OpenStudy (anonymous):

-2cos(pi) +2cos(0) -1/2cos(2pi) + 1/2cos(pi)

OpenStudy (xapproachesinfinity):

i didn't change the limits actually still 0 to pi

OpenStudy (anonymous):

i saw the wolfframalpha, ok but why did we plug in 2pi ?

OpenStudy (xapproachesinfinity):

well that's why i told you to not think about them jim just mentioned and error i made when i put the limits of the integral when i did U sub but since we went back to x again then we should not change the limits

OpenStudy (xapproachesinfinity):

not 2pi we plug is pi the function is cos 2x

OpenStudy (anonymous):

but i plugged in 2pi on -1/2cos(2pi)....

OpenStudy (xapproachesinfinity):

the integral gave us -1/2 cos(2x) remember?

OpenStudy (xapproachesinfinity):

so if we plug pi we get -1/2 cos (2 pi)

OpenStudy (anonymous):

ohhhh okay!!! i get it now!

OpenStudy (xapproachesinfinity):

Okay good!

OpenStudy (anonymous):

I got 3 now, I am getting close lol !

OpenStudy (xapproachesinfinity):

okay perhaps you forgot something there just check

OpenStudy (anonymous):

the -1/2 + 1/2 cancel out ....so I should get 4, an error in my signs

OpenStudy (xapproachesinfinity):

yup exactly :)

OpenStudy (xapproachesinfinity):

good job